OpenStudy (anonymous):

Module 1- Challenge Question 1 Hi guys, I was going through the questions for module 1 and I'm not sure how they got the relations (1.4) from equation 1.3. Any elaboration would be helpful, thank you.

OpenStudy (anonymous):

specifically how they got the relations 0 = -3V +W + X + 2(Y + Z) 0 = V 1 = -2X from [T] = [M*L ^{-3}]^{V} [L]^{W}[L*T ^{-2}]^{X}[L ^{2}]^{Y}[L ^{2}]^{Z}\]

OpenStudy (unklerhaukus):

\[\qquad\qquad [T] = [M\cdot L ^{-3}]^{V} [L]^{W}[L\cdot T ^{-2}]^{X}[L ^{2}]^{Y}[L ^{2}]^{Z}\\ ~\\ [M]^0[L]^0[T]^1 =[M]^V\cdot[L]^{-3V+W+X+2Y+2Z}[T]^{-2X}\] the dimensions are independent so \[ [M]^0=[M]^V\qquad\implies\qquad 0=V\\~\\ [T]^1=[T]^{-2X}\qquad\implies\qquad1=-2X\\\hspace{15em}X=-1/2\\~\\ [L]^0=[L]^{-3V+W+X+2Y+2Z}\qquad\implies\qquad 0=-3V+W+X+2Y+2Z\\\hspace{20em}0=0+W-1/2+2Y+2Z~\\\]